OpenStudy (anonymous):

Derek distributed eighteen pieces of apple to his friends. If each apple slice was one-eighth of an apple, how many apples did he distribute to his friends?

OpenStudy (anonymous):

18 * 1/8 = 18/8 = 2 1/4 apples